in an inter-barangay basketball game league, the team from hermana fausta has won 12 out of 25 games, a winning percentage of 48%. we have seen that they need to win 8 games consecutively to raise their percentage to 60%. what will their winning percentage if they win 10,20,30,50,100 games? can they reach a 100% winning percentage?
1
Expert's answer
2020-12-13T18:58:35-0500

Let n = number of wins

and p = number of games played

Winning percentage = n/p x 100%

When 10 games are won, n = 10 + 12 = 22 and p = 25 + 10 = 35

Winning percentage = 22/35 x 100% = 62.9%

When 20 games are won, n = 20 + 12 = 32 and p = 25 + 20 = 45

Winning percentage = 32/45 x 100% = 71.1%

When 30 games are won, n = 30 + 12 = 42 and p = 25 + 30 = 55

Winning percentage = 42/55 x 100% = 76.34%

When 50 games are won, n = 50 + 12 = 62 and p = 25 + 50 = 75

Winning percentage = 62/75 x 100% = 82.67%

When 100 games are won, n = 100 + 12 = 112 and p = 25 + 100 = 125

Winning percentage = 112/125 x 100% = 89.6%

The winning percentage increases as the number of games won increases, it's not possible to reach a 100% winning percentage because they already lost some games, they can only get closer as the number of games won increases.